# YSO Jets Driven by Magnetic Pressure Generated through Stellar Magnetosphere–Disk Interaction

### v1

@@ 5.1 Magnetospheric accretion and outflow
Figure 12: Accretion and outflow properties. Panel (a) shows a 3D rendition of the accretion streams by only choosing cells with real $\rho>10^{-13}~{}\mathrm{g~{}cm^{-3}}$ and kinetic-$\beta<1$. Panel (b) quantifies the landing locations of the mass accretion rate streams on the star at each polar angle, showing off-midplane accretion along magnetosphere fieldlines. Panel (c) shows the total mass accretion rate (black line on top), jet outflow rate (red line at the bottom), and the disk wind outflow rate (blue line in the middle). The red dashed line is the jet outflow rate using the real gas only, showing the density floor’s minimal effect on the outflow. Panel (d) shows the “Keplerian angular momentum radius” ($R_{a}$, equ. 29) of the jet (lower red line) and the disk wind (upper blue line). The horizontal dashed lines mark the averaged $R_{a}$ over the time shown.
5.2 Comparing to the disk-field-only jet model
@@
An alternative jet-launching mechanism involves jets driven solely by the disk magnetic field. Y. Tu et al. (2025b) and Y. Tu et al. (2025a) model outflows in the absence of a stellar magnetosphere and identify a mechanism in which jets are powered by the so-called “avalanche-accretion stream” in the magnetically elevated atmosphere of the disk. In this section, we compare the outflows driven by stellar magnetosphere-disk interactions with those produced by the disk magnetic field alone.
@@
To set the stage for this comparison, we first summarize the disk-field-only jet-launching mechanism described in Y. Tu et al. (2025b) and Y. Tu et al. (2025a). In this scenario, the disk is threaded by a large-scale, open vertical magnetic field. Due to differential rotation and magnetic field geometry, the azimuthal magnetic field $B_{\phi}$ must switch sign somewhere between the upper and lower hemispheres, typically along the disk surface on one side. This sign reversal leads to strong azimuthal magnetic braking, which removes angular momentum and drives an avalanche-accretion stream along the disk surface. The differential rotation between this stream and the open field lines rapidly amplifies the toroidal magnetic field, generating strong toroidal magnetic pressure that lifts a portion of the accretion stream vertically, launching a jet. A similar process drives a slower, more mass-loaded disk wind at larger cylindrical radii, though this does not require the formation of an avalanche-accretion stream. Some regions of the inner disk wind can be magnetorotational instability (MRI)-active, where turbulent gas is lifted to high altitudes before merging with the disk wind.
@@
On the side of the disk without the azimuthal field sign change, the avalanche-accretion stream does not form, but jets can still emerge if the polar region remains relatively mass-free. However, MRI-active turbulent disk winds can fill the polar region with turbulent gas in many cases, overloading the outflow and effectively choking the jet. As a result, disk-field-only models frequently produce one-sided jets (see also, e.g., W. Béthune et al., 2017; D. Takaishi et al., 2024).
A key distinction between the disk-only model and our magnetosphere-disk model is the persistence of a bipolar jet in our simulations (see Figure 1). This difference arises from the global magnetic field geometry. In the presence of a stellar magnetosphere, the opening of the outer magnetosphere establishes a poloidal magnetic field structure with opposite directions between the polar region and its immediate surroundings. With differential rotation between gas at different altitudes, this configuration produces a quadripartite global toroidal magnetic field topology, with bipartite in each hemisphere in the case of a rapidly spinning stellar magnetosphere (as shown in Figure 13). Magnetic reconnection occurs along the boundary between the two magnetic partitions in each hemisphere, naturally limiting the disk’s vertical expansion. This effect is analogous to the role of avalanche accretion streams in the disk-only models. As a result, the magnetic configuration inherently supports the formation of a bipolar jet. In contrast, the disk-field-only model maintains a bipartite global toroidal field geometry (see, e.g., Figure 9 in Y. Tu et al., 2025a), meaning that the toroidal field remains unipartite in each hemisphere. As a result, the avalanche accretion stream can develop only in the hemisphere where the toroidal field reverses direction. This asymmetry leads disk-field-only models to preferentially launch unipolar jets.
Figure 13: Illustration of the resulting large-scale toroidal magnetic field configuration, shaped by the poloidal field geometry established through the presence and partial opening of the rotating stellar magnetosphere. In contrast to the disk-only case (Y. Tu et al., 2025a), where the magnetic field direction reverses in only one hemisphere, both the poloidal and toroidal components exhibit reversal in both the upper and lower hemispheres in the case with a stellar magnetosphere.
@@
In contrast, the disk-magnetosphere model produces the toroidal field in an opposite manner. Here, strong magnetic field lines anchored to the rotating star transfer angular momentum to the disk surface, spinning up the upper layers of the disk relative to the denser regions below. As a result, a toroidal field is generated by the differential rotation between the faster-rotating upper layers and slower-rotating lower layers of the disk.
There are several important consequences of these differences in the generation of toroidal magnetic fields. One immediate effect is on the geometry of the poloidal magnetic field that channels the jet. In the disk-field-only model, the avalanche-accretion stream drags magnetic field lines inward, resulting in an inward-pinched poloidal field configuration (see Figure 5 in Y. Tu et al. 2025b). In contrast, in the magnetosphere-disk model, the pinching point of the field moves outward, leading to an outward-pinched poloidal magnetic field structure (Figure 7).
The difference in magnetic field geometry stems from differences in the power source of the jet. In the disk-field-only model, the disk itself provides all the energy and angular momentum for the outflow. In contrast, in the magnetosphere-disk model, the field lines anchored on the rotating star provide an additional source of energy and angular momentum to power the jet. In addition, once opened, the magnetosphere supplies the inner disk with a much stronger open poloidal field than in the disk-field-only case, and provides a global field configuration with oppositely directed open poloidal fields near the (elevated) disk surface that is conducive to magnetic reconnection and jet mass loading. As a result, jets in the magnetosphere-disk model are both more energetic and more massive than those in the disk-field-only model from Y. Tu et al. (2025a). Specifically, the jet in our simulations reaches velocities of several hundred km s-1, compared to $\approx 100$ km s-1 in the disk-field only case. The mass outflow rate is also higher, with $\dot{M}_{\mathrm{jet}}\approx 2\times 10^{-9}~{}M_{\odot}\ \mathrm{yr}^{-1}$ compared to $\approx 2\times 10^{-11}~{}M_{\odot}\ \mathrm{yr}^{-1}$. Additionally, the jet carries more angular momentum, with the Keplerian angular momentum radius (equ. 29) of $R_{a}\approx 0.6~{}\mathrm{au}$, compared to $0.3~{}\mathrm{au}$ in the disk-field-only model.
@@
5.3 Comparison with previous 2D models
The “load-fire-reload” jet-launching mechanism in our model shares many similarities with previous 2D models. Studies such as A. P. Goodson & R. M. Winglee (1999); T. Kuwabara et al. (2000); C. Fendt & D. Elstner (2000); S. Matt et al. (2002); M. M. Romanova et al. (2005); C. Fendt et al. (2011); M. Čemeljić et al. (2013); C. Zanni & J. Ferreira (2013) also observed the cyclic behavior of magnetic field lines, where outflows are launched before the field resets to its initial configuration for the next cycle. The formation of a current sheet due to differential rotation, as described in Sec. 4.2, is conceptually similar to the current sheet discussed in D. A. Uzdensky (2002).
However, our model differs from these 2D studies in two key ways: (1) the inclusion of the azimuthal dimension, and (2) the longer simulation runtime, allowing the system to reach a quasi-steady state. The additional dimension enables axisymmetric processes to break down, while the extended simulation duration ensures that transient effects from the initial conditions dissipate. These differences lead to three main consequences.
First, the jet can be launched continuously (see Figure 1) because magnetic expansion and reconnection can occur simultaneously at different azimuthal angles. This sustained flaring and reconnection stabilize the system into a quasi-steady state in the late simulation stages. Additionally, mass loading onto the jet-launching field lines becomes more efficient due to reconnection along the disk surface-polar cavity boundary. The frequent localized magnetic reconnection and field expansion events alleviate the need for the global large-scale episodic outbursts often seen in 2D simulations.
Second, the magnetic field geometry associated with jet launching differs from 2D models. In our simulation, the field does not need to fully reconnect into closed loops across the disk midplane before another jet-launching event can occur: the only requirement is that the opened magnetosphere field lines threading the star touch and at least shallowly embedded in the disk atmosphere (as examplified by the green and magenta field lines in Fig. 7); the touch of the disk is facilitated by the elevation of the atmosphere of the disk, driven by the amplification of the toroidal field in the inner ideal MHD part of the disk (J. S. Oishi et al., 2020). As a result, only localized reconnection along the surface of the highly elevated and highly dynamic disk atmosphere is necessary to generate a jet, rather than global reconnection involving a coherent poloidal field threading the disk midplane.
These differences, particularly the suppression of large-scale temporal variations, are consistent with the findings of S. Takasao et al. (2022). While 2D models are valuable for developing first-step theories and characterizing general properties, these differences highlight the crucial role of 3D models in capturing additional key ingredients of the outflow-launching process.
5.4 Connection to the axisymmetric steady-state outflow
Although the outflow launching mechanism in our model is highly variable in space and time, it is still valuable to put our model in the context of the axisymmetric steady-state wind theory.
Four quantities are conserved along a magnetic field line in the axisymmetric steady-state wind model. However, since our jet-launching mechanism (Sec. 4) involves the opening and reconnection of magnetic field lines, this “along a magnetic field line” picture is not strictly applicable. Nevertheless, we can reinterpret these conserved quantities to gain insight into the properties of the outflow.
The first conserved quantity,
$k=\frac{4\pi\rho v_{\mathrm{pol}}}{B_{\mathrm{pol}}},$ (30)
characterizes the degree of mass loading of the outflow. Physically, $k$ is the $4\pi$ times the ratio between the outflow mass flux $\dot{M}_{\mathrm{out}}=\rho v_{\mathrm{pol}}A$ and the magnetic flux $\Phi_{B}=B_{\mathrm{pol}}A$, where $A$ is an area element perpendicular to the poloidal field.
The second is the conserved angular velocity:
$\omega_{s}=\frac{v_{\phi}}{R}-\frac{kB_{\phi}}{4\pi\rho R},$ (31)
where the first term is the angular velocity of the gas and the second term can be interpreted as the magnetic contribution to the conserved angular velocity.
Similarly, the third quantity is the angular momentum, and the two terms are the angular momenta carried by the gas and magnetic field, respectively:
$l=Rv_{\phi}-\frac{RB_{\phi}}{k}.$ (32)
The last is the energy
$E=\frac{1}{2}v^{2}+\Phi-\frac{\omega_{s}RB_{\phi}}{k}+h,$ (33)
which is derived from Bernoulli’s equation. The four terms are energy carried by gas advection, gravity, magnetic field, and gas internal energy, respectively. Since thermal energy is negligible in the outflow (plasma-$\beta\ll 1$, Figure 4), the energy carried by the gas is essentially its kinetic energy.
To connect our model with these quantities, we show in Figure 14 the mass-weighted averaged quantities at every height $z$ in the jet, defined as gas moving at a speed $v_{z}^{p}>10^{7}~{}\mathrm{cm~{}s^{-1}}$. Note that while this analysis does not trace along individual magnetic field lines, the mass-weighted averaged quantities offer insight into the jet properties.
Figure 14(a) shows $k$, which is roughly constant along the jet. This indicates that the mass loading of the outflow remains approximately the same as the jet propagates outward.
@@ 5.4 Connection to the axisymmetric steady-state outflow
Figure 14(b), (c), and (d) show $\omega_{s}$, $l$, and $E$, and their gas and magnetic component, respectively. While the magnetic field dominates $\omega_{s}$ (the magnetic field is mostly toroidal) at large distances, gas dominates both angular momentum $l$ and energy $E$. Only around the base of the jet $\lesssim 2~{}\mathrm{au}$ does the magnetic field dominate in angular momentum and energy, suggesting that energy and angular momentum are transferred from the magnetic field to the gas as the outflow propagates outward, and the outflow on the large scale is gas-dominated. These findings are consistent with, and similar to, the analysis along the field line in Y. Tu et al. (2025b, a); J. Jacquemin-Ide et al. (2021).
To further quantify the outflow properties, we show in Figure 14(e) the plasma-$\beta$, kinetic-$\beta$ ($\beta_{K}$), and the poloidal kinetic-$\beta$, defined as
$\beta_{K,\mathrm{pol}}=\frac{\rho v_{\mathrm{pol}}^{2}}{B_{\mathrm{pol}}^{2}/4\pi}.$ (34)
While plasma-$\beta$ is confirmed to be much less than order unity, kinetic-$\beta$ stays above $1$ beyond the launching zone, emphasizing the gas dominance at large heights. The $\beta_{K,\mathrm{pol}}$ is larger than $\beta_{K}$, highlighting the dominance of the poloidal gas motion over the poloidal magnetic field. Figure 14(f) further quantifies the velocity and magnetic field geometry, showing the ratio between the toroidal and the poloidal velocity or magnetic field. At a large height, the poloidal velocity dominates the toroidal velocity, but the magnetic field is mostly toroidal. This is consistent with the observation in Figure 1(3D figure), where the streamlines (black lines) are mainly poloidal, while the magnetic fields (white) are mainly toroidal on the plotted scales.
Figure 14: Mass-weighted average conserved quantities (upper four panels) and outflow properties (lower two panels) at different heights. Panels (a), (b), (c), and (d) show the mass loading parameter (equ. 30), angular velocity (equ. 31), angular momentum (equ. 32), and energy (equ. 33) respectively. Panel (e) shows the plasma-$\beta$ (equ. 12), kinetic-$\beta$ (equ. 13), and the poloidal kinetic-$\beta$ (equ. 34). Panel (f) shows the ratios of toroidal to poloidal components for both velocity and magnetic field, illustrating the changing dominance of flow and field geometry with height.
6 Conclusions
We performed a 3D non-ideal MHD simulation of outflow-launching involving a rotating protostar, the stellar magnetosphere, the circumstellar disk, and open magnetic field lines threading the disk. In our model, a fast, collimated bipolar jet is launched, surrounded by a slower, wider, but more mass-loaded disk wind. Our main conclusions are as follows.

### v2

@@ V.1 Magnetospheric accretion and outflow
Figure 12: Accretion and outflow properties. Panel (a) quantifies the landing locations of the mass accretion rate streams on the star at each polar angle, showing off-midplane accretion along magnetosphere fieldlines. Panel (b) shows the total mass accretion rate (black line on top), jet outflow rate (red line at the bottom), and the disk wind outflow rate (blue line in the middle). The red dashed line is the jet outflow rate using the real gas only, showing the density floor’s minimal effect on the outflow. Panel (c) shows the “Keplerian angular momentum radius” ($R_{a}$, equ. 29) of the jet (lower red line) and the disk wind (upper blue line). The horizontal dashed lines mark the averaged $R_{a}$ over the time shown.
V.2 Comparing to the disk-field-only jet model
@@
An alternative jet-launching mechanism involves jets driven solely by the disk magnetic field. Tu et al. (2025b) and Tu et al. (2025a) model outflows in the absence of a stellar magnetosphere and identify a mechanism in which jets are powered by the so-called “avalanche-accretion stream” in the magnetically elevated atmosphere of the disk. In this section, we compare the outflows driven by stellar magnetosphere-disk interactions with those produced by the disk magnetic field alone.
@@
To set the stage for this comparison, we first summarize the disk-field-only jet-launching mechanism described in Jacquemin-Ide et al. (2021), Tu et al. (2025b), Tu et al. (2025a), and Mori et al. (2025). In this scenario, the disk is threaded by a large-scale, open vertical magnetic field. Due to differential rotation and magnetic field geometry, the azimuthal magnetic field $B_{\phi}$ must switch sign somewhere between the upper and lower hemispheres, typically along the disk surface on one side. This sign reversal leads to strong azimuthal magnetic braking, which removes angular momentum and drives an avalanche-accretion stream along the disk surface. The differential rotation between this stream and the open field lines rapidly amplifies the toroidal magnetic field, generating strong toroidal magnetic pressure that lifts a portion of the accretion stream vertically, launching a jet. A similar process drives a slower, more mass-loaded disk wind at larger cylindrical radii, although this does not require the formation of an avalanche-accretion stream. Some regions of the inner disk wind can be magnetorotational instability (MRI)-active, where turbulent gas is lifted to high altitudes before merging with the disk wind.
@@
On the side of the disk without the azimuthal field sign change, the avalanche-accretion stream does not form, but jets can still emerge if the polar region remains relatively mass-free. However, MRI-active turbulent disk winds can fill the polar region with turbulent gas in many cases, overloading the outflow and effectively choking the jet. As a result, disk-field-only models frequently produce one-sided jets (see also, e.g., Béthune et al., 2017; Takaishi et al., 2024).
A key distinction between the disk-only model and our magnetosphere-disk model is the persistence of a bipolar jet in our simulations (see Figure 1). This difference arises from the global magnetic field geometry. The partial opening of the stellar magnetosphere introduces a substantial amount of open flux that threads the star in the polar regions. This open magnetic flux results in a strong field within the polar cavity, one that dominates over the disk magnetic field. As a result, the low-density polar cavity remains open in both hemispheres, allowing a stable bipolar jet to develop. In contrast, the disk-field-only model lacks this concentrated polar magnetic flux. Without the stellar contribution, the low-density jet cavity is more easily filled by the high-density disk-wind material, especially on the side where avalanche accretion does not occur. This configuration naturally leads to an asymmetric jet structure and, in some cases, a strongly one-sided jet.
@@
In contrast, the disk-magnetosphere model produces the toroidal field in an opposite manner. Here, strong magnetic field lines anchored to the rotating star transfer angular momentum to the (elevated) disk surface, spinning up the upper layers of the disk relative to the denser regions below. As a result, a toroidal field is generated by the differential rotation between the faster-rotating upper layers and slower-rotating lower layers of the disk.
There are several important consequences of these differences in the generation of toroidal magnetic fields. One immediate effect is on the geometry of the poloidal magnetic field that channels the jet. In the disk-field-only model, the avalanche-accretion stream drags magnetic field lines inward, resulting in an inward-pinched poloidal field configuration (see Figure 5 in Tu et al. 2025b). In contrast, in the magnetosphere-disk model, the pinching point of the field moves outward, leading to an outward-pinched poloidal magnetic field structure (Figure 7).
The difference in magnetic field geometry stems from differences in the power source of the jet. In the disk-field-only model, the disk itself provides all the energy and angular momentum for the outflow. In contrast, in the magnetosphere-disk model, the field lines anchored on the rotating star provide an additional source of energy and angular momentum to power the jet. In addition, once opened, the magnetosphere supplies the inner disk with a much stronger open poloidal field than in the disk-field-only case, and provides a global field configuration with oppositely directed open poloidal fields near the (elevated) disk surface that is conducive to magnetic reconnection and jet mass loading. As a result, jets in the magnetosphere-disk model are both more energetic and more massive than those in the disk-field-only model from Tu et al. (2025a). Specifically, the jet in our simulations reaches velocities of several hundred km s-1, compared to $\approx 100$ km s-1 in the disk-field only case. The mass outflow rate is also higher, with $\dot{M}_{\mathrm{jet}}\approx 2\times 10^{-9}~M_{\odot}\ \mathrm{yr}^{-1}$ compared to $\approx 2\times 10^{-11}~M_{\odot}\ \mathrm{yr}^{-1}$. Additionally, the jet carries more angular momentum, with the Keplerian angular momentum radius (equ. 29) of $R_{a}\approx 0.6~\mathrm{au}$, compared to $0.3~\mathrm{au}$ in the disk-field-only model.
@@
V.3 Magnetic-to-Kinetic Energy and Angular Momentum Conversion in the Jet
There are four conserved quantities along magnetic field lines in the axisymmetric steady-state wind model (see, e.g., equation 32 to 35 in Tu et al. 2025a; also Salmeron et al. 2011; Zhu and Stone 2018; Jacquemin-Ide et al. 2021). However, since our jet-launching mechanism (Sec. IV) is highly variable, involving repeated opening and reconnection of field lines, these quantities are not strictly conserved along individual field lines. Nevertheless, we can still make a rough connection to three of them using the total fluxes of mass, poloidal field, angular momentum, and energy integrated over the cross-sectional area of the jet at different heights as follows.
The first quantity
$k^{\prime}=\frac{\int_{\mathrm{jet}}4\pi\rho v_{\mathrm{z}}~dA}{\int_{\mathrm{jet}}B_{\mathrm{z}}~dA}$ (30)
is the ratio of the mass flux to the poloidal magnetic flux threading the jet. The second quantity
$l^{\prime}=\frac{\int_{\mathrm{jet}}\big[\rho v_{\mathrm{z}}(Rv_{\phi})-\frac{RB_{\phi}B_{\mathrm{z}}}{4\pi}\big]~dA}{\int_{\mathrm{jet}}\rho v_{\mathrm{z}}dA}$ (31)
is the angular momentum flux of the jet per unit mass flux. The third quantity
$E^{\prime}=\frac{\int_{\mathrm{jet}}\big[\frac{1}{2}\rho v^{2}v_{\mathrm{z}}+\frac{B^{2}v_{\mathrm{z}}}{4\pi}-\frac{(\mathbf{B}\cdot\mathbf{v})B_{\mathrm{z}}}{4\pi}\big]dA}{\int_{\mathrm{jet}}\rho v_{\mathrm{z}}dA}$ (32)
is the total (kinetic and magnetic) energy flux per unit mass flux. Both thermal and gravitational terms are ignored since they are negligible compared to the kinetic and magnetic terms.
@@ V.3 Magnetic-to-Kinetic Energy and Angular Momentum Conversion in the Jet
Figure 13(a) shows $k^{\prime}$, which is roughly constant along the jet. This indicates that the mass loading of the outflow remains approximately the same as the jet propagates outward. Figure 13(b) and (c) show $l^{\prime}$, and $E^{\prime}$, and their gas and magnetic component, respectively. Near the jet base ($\lesssim 2~\mathrm{au}$), the magnetic field dominates the transport of both angular momentum and energy. Farther out, the angular momentum carried by the gas becomes comparable to that of the magnetic field, while the energy flux $E^{\prime}$ is increasingly dominated by the gas component. This transition suggests an ongoing transfer of energy and angular momentum from the magnetic field to the gas as the outflow expands, leading to a progressively gas-dominated jet at larger scales. The total angular momentum carried by the jet is substantial, especially when the magnetic contribution is included, which exceeds the gas-only angular momentum estimated in Sec. V.1. These findings are consistent with, and similar to, the analysis along the field line in Tu et al. (2025b, a); Jacquemin-Ide et al. (2021).
Figure 13: Integrated jet quantities defined in equations 30-32 as a function of height $z$: (a) the mass flux to magnetic flux ratio, (b) the angular momentum flux per unit mass flux, and (c) the energy flux per unit mass flux, showing the magnetic domination of the angular momentum and energy fluxes near the jet base and conversion of magnetic angular momentum and energy to the gas at larger distances.
V.4 Towards a Scenario of Accretion-fed Polar Poynting Jet Driven Along Open Stellar Field Lines Through Disk-Magnetosphere Interaction
Figure 14: Flow quantities averaged azimuthally and over time (from 3.5 to 4.5 years) to highlight the differences between the low-density magnetically dominated fast polar jet and its surrounding, denser, matter-dominated more slowly moving disk atmosphere/disk wind in (a) density, (b) vertical velocity away from the midplane, (c) the regular (thermal) plasma-$\beta$, and (d) the kinetic plasma-$\beta_{K}$ on a relatively small scale, and (e) vertical velocity away from the midplane on a larger scale. Magnetic field lines are plotted in each panel, showing that the low-density, magnetically dominated, fast-moving jet (with $v_{\mathrm{z}}^{p}>10^{7}$cm/s) flows outward primarily along open stellar field lines.
One of the most striking features of our simulation is that the fast, collimated jet material appears to flow primarily along open magnetic field lines that are anchored on the central star. It can already be seen from the azimuthally averaged flow quantities plotted in Fig. 7[b] (and its associated animation; see also Fig. 4[b] on a larger scale), and is illustrated more clearly in Fig. 14, where we show flow quantities averaged both azimuthally and over time (from 3.5 to 4.5 yrs or about 30 stellar spin periods).
The magnetic field configuration in Fig. 14 is similar to that in Fig. 1 of Zanni and Ferreira (2013), which displayed the time-averaged magnetic field and flow patterns from their reference 2D (axisymmetric) disk-magnetosphere interaction simulation with a prescribed magnetic diffusivity and fluid viscosity. They found four distinctive regions: (1) a stellar wind, where a fast outflow is driven thermally along open stellar field lines, (2) a disk wind, where a slower outflow is driven by field lines anchored on the disk, (3) a magnetospheric ejection (ME) region between the disk wind and stellar wind, where field lines connecting both the star and the disk are episodically inflated and reconnected, ejecting plasmoids with detached field lines, and (4) a region where the stellar magnetic field is more persistently connected to the disk.
Since our fast-moving jet primarily flows out along open stellar field lines, it would be called a stellar wind in Zanni and Ferreira’s terminology. However, there is a crucial difference between our jet and their stellar wind. The material in their stellar wind comes from the stellar surface, accelerated outwards by a prescribed thermal pressure along the open stellar field lines. In contrast, no hot gas capable of escaping to large distances is prescribed near the stellar surface in our simulation; as a result, the gas near the stellar surface is pulled by strong gravity inward toward the star along the open field lines in the polar region rather than flowing outward (see Fig. 14[b]). The outflowing material in our simulation appears to be loaded onto the open stellar field lines well above (and away from) the stellar surface, likely through magnetic reconnection, which is expected to be more chaotic and localized in our 3D global simulation without a prescribed magnetic diffusivity and fluid viscosity (as discussed below), although numerical diffusion may also contribute. Determining the exact origin of the jet mass loading is a pressing issue that needs to be addressed in future investigations, such as through higher resolution simulations to reduce numerical diffusion.
Although the jet material ultimately flows along open stellar magnetic field lines, these lines are not necessarily attached only to the star at all times. They typically undergo a cyclic evolution in which a given field line is first connected simultaneously to the star and the disk (or its elevated, turbulent atmosphere), allowing dense disk gas and magnetic stress to be transferred onto it. Magnetic reconnection then severs the disk footpoint, leaving the line open and anchored only on the star; the toroidal magnetic pressure built up during the connected phase is released, accelerating the newly loaded plasma through the low-density polar cavity (which is kept open by the star-anchored, predominantly open field lines in the cavity in the first place). In this way, the jet is instantaneously a star-anchored outflow, but is powered and mass-loaded by earlier episodes in which the now-open field lines interact with the denser disk/atmosphere. This transient two-footpoint phase is essential for understanding how a magnetically driven, star-anchored Poynting jet can remain continuously energized and supplied with mass.
The differences between our work and Zanni and Ferreira (2013) in the level of (explicit) magnetic diffusivity and in the simulation’s dimensionality have important consequences. First, the absence of (explicit) magnetic diffusion enables rapid winding of the initially weak magnetic field in the inner disk, leading to a buildup of magnetic pressure that forms a highly elevated, dynamic (indeed MRI-unstable) disk atmosphere, which gradually merges into a slow, dense disk-wind. The puff-up facilitates a more continuous jet driving because the open stellar field lines (which have a natural tendency to curve back toward the midplane and return to the initial dipolar configuration) only need to touch, or shallowly embedded in, the elevated disk atmosphere (as illustrated by the green and magenta field lines in Fig. 7) to drive outflow. These “touch-and-go” field lines, which are a subset of the “two-legged” field lines, increase the rapidness of mass ejection by not requiring a full reset of the global magnetic field geometry. Just as importantly, the extra dimension in our 3D simulation enables magnetic reconnection to occur asynchronously across azimuth—concurrently but without global coherence—so that different $\phi$ sectors load and release mass and magnetic energy independently. It makes the local reconnection-driven “load-fire-reload” jet-launching cycle much shorter than those observed in 2D simulations, which can last for several stellar rotation periods (e.g., Zanni and Ferreira, 2013). The more frequent and localized reconnection events enabled by the third (azimuthal) dimension and the disk puff-up alleviate the need for the global large-scale episodic outbursts often seen in 2D simulations. They facilitate the rather continuous, quasi-steady mass ejection seen in our jet (see the animated version of Fig. 1), particularly at later times after the initial adjustments. The suppression of large-scale temporal variations in 3D is consistent with the findings of Takasao et al. (2022).
Our polar jet can be viewed as a variant of the so-called “accretion-powered stellar wind” advocated by Matt and Pudritz (2005, compare our Fig. 14a,b with the cartoon sketched in their Fig. 1) in the context of braking the stellar spin. They envisioned several processes for transferring some of the accretion energy to the corona region of open stellar field lines, such as accretion shocks, magnetic reconnection, and magnetic waves, which may raise the stellar corona’s temperature and thermally drive a stellar wind, as simulated by Zanni and Ferreira (2013) and others (e.g., Matt and Pudritz, 2008). However, our jet has a thermal pressure orders of magnitude smaller than the magnetic pressure (see Fig. 14[c]) and is thus completely magnetically driven; it is essentially an electromagnetic (Poynting) jet powered by field line bending through the interaction between the stellar field lines and the surrounding disk (and its elevated atmosphere and wind), with the Poynting energy flux gradually converted into the kinetic energy flux as the jet propagates outward (see Fig. 13[c]).
The Poynting jet in our simulation has more in common with the magnetically dominated axial jet found in the 2D (axisymmetric) simulations of Ustyugova et al. (2006, see also ) in the strong “propeller” regime of disk-magnetosphere interaction than with the thermally driven stellar wind. One major difference is that they prescribed an explicit magnetic diffusivity and fluid viscosity, as in Zanni and Ferreira (2013). The explicit magnetic diffusivity allows the material flowing along the upper funnel flow region of the closed magnetosphere to diffuse to its neighboring open stellar field lines (see, e.g., Fig. 11d of Ustyugova et al. 2006 or Fig. 19 of Romanova et al. 2009), thereby providing an elegant explanation for the jet mass loading. Our 3D simulation shows that mass loading onto open stellar field lines is possible even without explicit magnetic diffusivity, likely due to a more chaotic flow in both space and time, as mentioned earlier, which results in an effective cross-field mass diffusion, plausibly between the open stellar field and both the closed magnetospheric field carrying funnel flows and the field threading the elevated disk atmosphere/disk-wind bordering the jet.
The explicit magnetic diffusivity and viscosity in Ustyugova et al. (2006) is also likely the reason why their rather laminar accretion flow can cross well-ordered poloidal field lines near the inner edge of the disk and be ejected as a conic wind surrounding the axial jet (see their Fig. 11). In contrast, the absence of an explicit magnetic diffusivity in our simulation facilitates field line winding in and around the inner disk through differential rotation, which puffs up the disk atmosphere and drives a broader disk-wind surrounding the axial jet (see Fig. 11[a]).
Another difference is that Ustyugova et al. (2006) only found strong jets in the strong “propeller” regime. In our simulation, the truncation radius is comparable but slightly smaller than the corotation radius, placing the system in the non-propeller or, at most, weak propeller regime. However, a strong jet is still produced. One plausible way to reconcile this apparent contradiction is that the stellar field lines in our simulation are connected to the elevated disk atmosphere at higher heights and larger radii, where the gas rotates more slowly than at the inner disk edge on the midplane, creating an “effective propeller” regime above and below the disk that enables energy and angular momentum to be extracted magnetically from the rotating star to power the jet.
Nevertheless, in both our simulation and those of Ustyugova et al. (2006), a fast collimated magnetically dominated axial jet is produced despite significant differences in magnetic diffusivity, fluid viscosity, simulation dimensionality (2D vs 3D), and whether the system is in the strong “propeller” regime. This consistency in outcome points to a robust mechanism for producing the accretion-fed, tenuous axial Poynting jet driven in the low-density polar cavity along open stellar magnetic field lines by field line bending through the magnetosphere-disk interaction. Crucially, the interaction-opened stellar magnetic field pushes against the elevated disk atmosphere/disk wind to keep the low-density polar cavity open, enabling the light mass-loading critical for the jet acceleration to a high speed.
This lightly loaded, magnetically dominated, axial jet is physically distinct from the traditional disk wind launched over a range of disk radii or the X-wind launched near the inner edge of the disk. In particular, it is not the densest part of a magneto-centrifugal disk- or X-wind that extends all the way to the axis. A disk- and/or X-wind can still exist outside the axial Poynting jet. They would have the same general source of mass loading (accreting material either in the disk/atmosphere or funnel flow), but they would be launched by field lines anchored primarily on the disk rather than the star. As such, unlike the disk- or X-wind, the jet in our simulation can extract angular momentum from the star directly, although we refrain from a quantitative analysis of the stellar braking by the jet because of the uncertainties about the exact details of its mass loading.
It is possible that the part of the disk- and/or X-wind immediately surrounding the axial Poynting jet can be collimated toward the axis, forming an outer, more mass-loaded, layer of the global jet system at large distances. Furthermore, it is likely that the central YSO blows a tenuous (Sun-like) stellar wind, which may form yet another layer (likely the innermost and fastest) of the jet (see Ferreira et al. 2006 for a discussion) [^6] This stellar wind would likely prevent the infall in the polar region in our simulation (see, e.g., the left panel of Fig. 2), although its effects on the global 3D flow structure remain to be quantified.. The existence of different jet components is hinted at, for example, by the detection of an X-ray jet in DG Tau, which indicates a faster speed than measured in the optical/IR component of the jet (as sketched in, e.g., Fig. 3 of Günther et al. 2009. In any case, the existence of a Poynting jet as described in this work and that of Usytugova et al. does not preclude other jet components of different origins. Future investigations should focus on the relative importance of these potential components under different conditions and on their possible interactions.
VI Conclusions
We have carried out global, three-dimensional magnetohydrodynamic simulations of the interaction between a rotating, magnetized young star and its accretion disk, focusing on how the coupled system launches and sustains an axial jet. The simulation follows the evolution of both accretion and outflow without prescribing magnetic diffusivity or viscosity, allowing magnetic reconnection, mass loading, and jet launching to emerge more self-consistently. Our main findings are summarized below.